# PRISM [Uncaptioned image] of Opinions: A Persona‑Reasoned Multimodal Framework for User-centric Conversational Stance Detection

### v1

@@ Abstract
The rapid proliferation of multimodal social media content has driven research in Multimodal Conversational Stance Detection (MCSD), which aims to interpret users’ attitudes toward specific targets within complex discussions. However, existing studies remain limited by: 1) pseudo-multimodality, where visual cues appear only in source posts while comments are treated as text-only, misaligning with real-world multimodal interactions; and 2) user homogeneity, where diverse users are treated uniformly, neglecting personal traits that shape stance expression. To address these issues, we introduce U-MStance, the first user‑centric MCSD dataset, containing over 40k annotated comments across six real-world targets. We further propose PRISM, a Persona-Reasoned multImodal Stance Model for MCSD. PRISM first derives longitudinal user personas from historical posts and comments to capture individual traits, then aligns textual and visual cues within conversational context via Chain‑of‑Thought to bridge semantic and pragmatic gaps across modalities. Finally, a mutual task reinforcement mechanism is employed to jointly optimize stance detection and stance-aware response generation for bidirectional knowledge transfer. Experiments on U‑MStance demonstrate that PRISM yields significant gains over strong baselines, underscoring the effectiveness of user-centric and context-grounded multimodal reasoning for realistic stance understanding.
@@ 1 Introduction
Social media have become a dominant platform for the expression of opinions, making their analysis crucial to understand divisive societal topics such as politics, public health, and social justice [26-glandt2021stance, 27-liang2022zero]. Stance detection, the task of identifying a user’s position toward a target, has become vital for charting public discourse and polarization. As online interactions grow more complex, the field has evolved from simple textual analysis to modeling rich, multi-party conversational contexts [28-li2023improved, 29-li2023contextual, 30-niu2024challenge].
@@
The recent shift towards Multimodal Conversational Stance Detection (MCSD), exemplified by the MmMtCSD dataset [1-niu2024multimodal], marks a crucial step forward. However, despite this progress, two fundamental limitations persist across the field, hindering authentic stance modeling. The first is a pseudo-multimodality, where visual cues appear only in source posts while comments remain text‑only, diverging from the multimodal nature of real-world interactions. The second is user homogeneity, where diverse individuals are modeled uniformly, ignoring personal traits that shape stance expression. These limitations can be clearly observed in the example shown in Figure 1. User Carrots combines an image and textual content to contrast Obama and Trump, conveying a derogatory attitude toward Trump, which highlights the necessity of multimodal cues for accurate stance understanding. Meanwhile, Zeke expresses sympathy for Trump, and Bgad challenges this view, revealing conflicting stances. A model that remains blind to their individual personas would fail to grasp the reason for their disagreement, leading to incorrect stance interpretations.
To catalyze progress beyond these limitations, we therefore introduce U-MStance, the first dataset designed for user-centric multimodal conversational stance detection. Comprising over 40k annotated comments across six real-world targets, U-MStance uniquely incorporates multimodality in both source posts and conversational replies, providing a realistic and challenging benchmark that captures the nuances of user-centric, multimodal interactions. Building upon this critical new dataset, we further propose PRISM, a Persona-Reasoned multImodal Stance Model for MCSD. Specifically, PRISM first explores the underlying personal tendencies behind stance expressions and derives longitudinal user persona traits inspired by Big Five (OCEAN) theory [44-jiang2024personallm] from historical posts and comments. Then, it introduces a rationalized cross-modal grounding module, leveraging Chain-of-Thought (CoT) reasoning to explicitly bridge the profound semantic and pragmatic gap between modalities in conversational contexts. Ultimately, these personas and contextual representations are fed into a mutual task reinforcement mechanism, which jointly optimizes stance classification and stance-aware response generation, enabling bidirectional knowledge transfer that enhances both detection performance and the modeling of user communication styles. Comprehensive experiments on our U-MStance dataset show that PRISM achieves significant improvements over strong baselines across multiple metrics. The main contributions can be summarized as follows:
@@ 2.1 Stance Datasets
The development of stance detection datasets has evolved from analyzing isolated texts to modeling complex interactions. Early benchmarks like SemEval-2016 Task 6 [2-mohammad2016semeval] and VAST [3-allaway2020zero] focused on sentence-level textual stance detection, establishing foundational models but overlooking the context of dialogue. This limitation spurred the creation of conversational datasets such as SRQ [4-villa2020stance] and CANT-CSD [5-li2023improved], which treated stance as a dynamic process unfolding within a conversation. The most recent advancement is the shift toward multimodal stance detection, which incorporates visual elements alongside text to better reflect modern social media discourse. Despite these advances, even the latest datasets for MCSD, such as MmMtCSD [1-niu2024multimodal], still face two key shortcomings that hinder realistic stance modeling. First, multimodality is handled only superficially, with source posts treated as multimodal but conversational replies reduced to text-only. In addition, information about users, including their posting history, is largely ignored, overlooking contextual factors that shape an individual’s expressed stance and leaving a substantial gap in building comprehensive and authentic stance detection models.
@@ 2.2 Stance Detection Methods
Methodological advancements in stance detection has spanned multiple task settings in-target [6-li2021multi], cross-target [7-ding2024cross, 8-liang2021target, 9-wei2019modeling], and zero-shot scenarios [10-allaway2020zero, 11-liang2022zero, 12-zhu2022enhancing]. Early deep learning approaches primarily fine-tuned pre-trained models such as BERT [13-cambria2018senticnet]. However, the field was revolutionized by the emergence of Large Language Models (LLMs), whose powerful semantic understanding and adaptability introduced new capabilities [14-cheng2024emotion, 15-cheng2024mips, 16-li2024mitigating]. The ability of LLMs to perform zero-shot stance prediction through simple prompting, in particular, dramatically improved the task’s accessibility and performance [18-ding2024edda].
Following this paradigm shift, LLM-centric research has rapidly expanded, moving beyond direct prompting or CoT reasoning toward increasingly sophisticated methodologies [21-zhang2023investigating, 22-zhang2022would]. These include fine-tuning with external knowledge [19-li2023stance], prompt-based multi-perspective reasoning [23-ding2024cross, 17-wang2025more], formalizing logic structures [24-zhang2024knowledge], and designing multi-agent collaborative systems [20-lan2024stance, 25-wang2024deem]. Despite this rapid progress, current approaches continue to treat users as a homogeneous group, overlooking the personal traits and discourse tendencies that shape stance expression. This simplification limits the model’s ability to interpret conflicting viewpoints, capture personalized reactions, and reason about stance in realistic multi-party interactions. Bridging this gap requires moving beyond content-centered modeling toward a genuinely user-centric perspective.
@@ 3 Dataset
This section outlines the construction and key attributes of our U-MStance dataset, comprising 40,003 annotated texts and images. As illustrated in Figure 2, our data construction pipeline ensures high data quality and relevance, involving data collection and data processing to ensure high data quality. Subsequently, data analysis is conducted to profile the dataset’s composition and validate its quality.
@@ 3.1 Data Construction
Dataset Collection. To capture realistic user interactions, we sourced conversational data from Reddit [^1] https://www.reddit.com, a large and diverse online forum, using the PRAW library. We began by identifying initial posts related to six high-engagement, real-world targets: Trump, Biden, Tesla, BMW, Costco, and Bitcoin. Posts were selected based on popularity metrics like upvotes and comment counts to ensure a sufficient volume of conversational replies for in-depth annotation. We also collected each user’s historical posts and comments to enrich their personal information.
@@
Pre-annotation. Following this, GPT-4o-mini was employed for automated pre-annotation, assigning a preliminary stance label (Favor, Against, or None) to each text relative to its target.
Human Verification. The automated labels underwent a rigorous human verification process by seven NLP researchers. Annotators were first qualified through two pilot rounds and an expert review. During the main annotation, each instance was independently reviewed by at least two annotators. Disagreements were resolved by a third senior annotator, and the final labels were determined through majority voting. Beyond stance correction, annotators also evaluated the contextual relevance of any associated images. This multi-annotator workflow ensures high label fidelity, incorporates diverse expertise, and reinforces the overall reliability of the dataset.
Table 1: The comparison between our U-MStance with other stance detection datasets. Comment: the number of comment texts; Conv.: support for the conversational context; A-Turns: average turns per conversation; Img: support for the image modality; Target: support for the target type(noun/claim); User: support for user information; -: not mentioned in the original paper.
Dataset Comment Conv. A-Turns Img Target User
SEM16 [2-mohammad2016semeval] 4,870 1.00 noun
P-stance [31-li2021p] 21,574 1.00 noun
COVID-19 [26-glandt2021stance] 6,133 1.00 noun
WT-WT [32-conforti2020will] 51,284 1.00 noun
VAST [3-allaway2020zero] 23,525 1.00 noun
SRQ [4-villa2020stance] 1,348 1.00 noun
CANT-CSD [5-li2023improved] 5,376 11.75 claim
CTSDT [33-li2023contextual] 53,861 - noun
MT-CSD [30-niu2024challenge] 15,876 4.82 noun
C-MTCSD [34-niu2025c] 24,264 3.13 noun
MT2-CSD [35-niu2025mt2] 24,457 4.41 claim
MMVax [37-weinzierl2023identification] 11,300 1.00 noun
MMSD [36-liang2024multi] 17,544 1.00 noun
MmMtCSD [1-niu2024multimodal] 21,340 3.46 claim
U-MStance(Ours) 40,003 4.10 noun
Figure 3: The distribution of stance categories in our U-MStance.
Figure 4: Overflow of our PRISM framework.
@@ 3.2 Data Analysis
Data Statistics. Table 1 and Figure 3 provide a statistical overview of U-MStance, contextualizing it among existing stance detection datasets. Our dataset comprises 40,003 annotated instances. Crucially, as highlighted in Table 1, U-MStance is the first dataset in this domain to systematically incorporate user information, offering a foundational resource to address the prevalent issue of user homogeneity. Figure 3 illustrates the distribution of instances across different labels. For model development, the dataset is partitioned into training, validation, and test sets using an 8:1:1 ratio, ensuring a balanced representation for comprehensive evaluation and analysis.
Quality Assessment. To ensure the annotation reliability, we measured inter-annotator agreement using Cohen’s Kappa [38-cohen1960coefficient], achieving an average score of 0.64, indicative of substantial agreement. This consistency was further reinforced through a rigorous post-annotation review, in which senior researchers audited a randomly selected subset of the annotated data to verify accuracy and resolve borderline cases, providing an additional layer of quality assurance.
@@ 4.1 Longitudinal User Persona Distillation
Stance is not an isolated behavior but an externalized reflection of a user’s underlying personal tendencies. Thus, each user is modeled as a persistent yet evolving entity whose stance expression is influenced by stable personality traits. For this purpose, we adopt the empirically validated Big Five (OCEAN) personality model [44-jiang2024personallm] from psychology. This model characterizes personality along five fundamental dimensions: Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism. To infer these traits for user $u_{N}$, we first aggregate their complete multimodal history, $H_{u_{N}}$, which comprises all of their past posts and comments.
@@
where each rating $r\in\{1,2,3,4,5\}$. The resulting profile $p_{u_{N}}$ is a concise, structured text string that serves as a crucial, personalized context vector. This profile is explicitly injected into the prompt for the final stance detection task, ensuring the model’s prediction is conditioned on the stable identity of the user in question.
@@ 4.2 Rationalized Cross-Modal Grounding
Images in online conversations are not merely decorative; they are communicative acts whose pragmatic meaning is determined by the conversational context. To decipher this context-dependent rhetorical intent, we propose the Rationalized Cross-Modal Grounding (RCMG) module, which generates an intent-aware caption for each image in the conversation. This module employs a two-stage reasoning process, inspired by Chain-of-Thought [45-wei2022chain], to bridge the semantic-pragmatic gap.
@@
The resulting caption $\hat{x}_{i}$ captures the rhetorical function of the image within the conversational context.
@@ 4.3 Mutual Task Reinforcement
To cultivate a deep, user-centric understanding of stance, PRISM adopts a multitask learning paradigm that jointly optimizes two inherently correlated tasks. The central intuition is that by training the model to generate a user’s response (the auxiliary task), it gains a richer understanding of the pragmatic and user-specific cues necessary to accurately predict the stance itself (the primary task). This mutual reinforcement acts as a strong regularizer, enhancing the robustness and prediction accuracy.
@@
The classification loss is the negative log-probability of generating the ground-truth label $y_{N}$.
@@ 5 Experiment
In this section, we present the baseline models evaluated on our proposed dataset, encompassing the experimental setup with implementation details and evaluation metrics, a description of the baseline methods, and a summary of the experimental results achieved by our PRISM framework.
@@ 5.1 Experimental Setup
Implement details. Our framework is built upon the open-source Qwen2.5-VL-7B model [48-bai2025qwen2], and is fine-tuned end-to-end for 5 epochs using our multitask objective, with the loss-weighting hyperparameter $\lambda$ set to 0.7. Training is performed with the AdamW optimizer [51-adam2014method], an initial learning rate of $1\times 10^{-5}$, a cosine decay schedule, and a global batch size of 16 achieved through gradient accumulation. All experiments were run on a server with four NVIDIA A100 (40 GB) GPUs using PyTorch and the Hugging Face Transformers library.
Evaluation metrics. We adopt the average F1 score (F1-avg) as the primary evaluation metric, following [1-niu2024multimodal] and [50-niu2025c]. F1-avg is computed as the mean of F1 scores for the against and favor stances, denoted as F1-against and F1-favor. In addition, we also report the overall F1 score for comprehensive comparison.
@@ 5.2 Baseline Methods
We conduct extensive experiments with a range of existing methods, categorized into text-only and multimodal approaches. These baselines are evaluated across diverse stance detection scenarios, providing a thorough analysis of their effectiveness and highlighting the benefits of incorporating multimodal information into the stance detection.
Text-only baselines. We evaluate the following methods: (1) Pre-trained BERT [39-devlin2019bert], fine-tuned on the training data; (2) Pre-trained RoBERTa [40-liu2019roberta], an enhancement over BERT with larger batch sizes and more training data; (3) PoliBERTweet [43-kawintiranon2022polibertweet], a BERTweet-base model specialized for Twitter political posts. We also incorporate LLMs, including (4) LLaMA-2-7B (LLaMA2) [49-touvron2023llama], (5) GPT-3.5222https://platform.openai.com/docs/models/gpt-3-5 and (6) GPT-4333https://platform.openai.com/docs/models/gpt-4-turbo. Multimodal baselines. We include: (1) BERT-ViT [36-liang2024multi] which uses BERT for textual encoding and ViT for visual encoding; (2) TMPT [36-liang2024multi] employing targeted prompts for both pre-trained language and visual models; (3) LLaVA-1.5-7B (LLaVA) [47-liu2024improved]; (4) Qwen2.5-VL-7B (Qwen-VL) [48-bai2025qwen2]; (5) MiMo-VL-7B-SFT (MiMo) [46-xiaomi2025mimo]; and (6) GPT4-1444https://platform.openai.com/docs/models/gpt-4.1.

### v2

@@ Abstract
The rapid proliferation of multimodal social media content has driven research in Multimodal Conversational Stance Detection (MCSD), which aims to interpret users’ attitudes toward specific targets within complex discussions. However, existing studies remain limited by: 1) pseudo-multimodality, where visual cues appear only in source posts while comments are treated as text-only, misaligning with real-world multimodal interactions; and 2) user homogeneity, where diverse users are treated uniformly, neglecting personal traits that shape stance expression. To address these issues, we introduce U-MStance, the first user‑centric MCSD dataset, containing over 40k annotated comments across six real-world targets. We further propose PRISM, a Persona-Reasoned multImodal Stance Model for MCSD. PRISM first derives longitudinal user personas from historical posts and comments to capture individual traits, then aligns textual and visual cues within conversational context via Chain‑of‑Thought to bridge semantic and pragmatic gaps across modalities. Finally, a mutual task reinforcement mechanism is employed to jointly optimize stance detection and stance-aware response generation for bidirectional knowledge transfer. Experiments on U-MStance demonstrate that PRISM yields significant gains over strong baselines, underscoring the effectiveness of user-centric and context-grounded multimodal reasoning for realistic stance understanding.
@@ 1 Introduction
Social media have become a dominant platform for the expression of opinions, making their analysis crucial to understand divisive societal topics such as politics, public health, and social justice [13, 25]. Stance detection, the task of identifying a user’s position toward a target, has become vital for charting public discourse and polarization. As online interactions grow more complex, the field has evolved from simple textual analysis to modeling rich, multi-party conversational contexts [21, 23, 33].
@@
The recent shift towards Multimodal Conversational Stance Detection (MCSD), exemplified by the MmMtCSD dataset [31], marks a crucial step forward. However, despite this progress, two fundamental limitations persist across the field, hindering authentic stance modeling. The first is a pseudo-multimodality, where visual cues appear only in source posts while comments remain text‑only, diverging from the multimodal nature of real-world interactions. The second is user homogeneity, where diverse individuals are modeled uniformly, ignoring personal traits that shape stance expression. These limitations can be clearly observed in the example shown in Figure 1. User Carrots combines an image and textual content to contrast Obama and Trump, conveying a derogatory attitude toward Trump, which highlights the necessity of multimodal cues for accurate stance understanding. Meanwhile, Zeke expresses sympathy for Trump, and Bgad challenges this view, revealing conflicting stances. A model that remains blind to their individual personas would fail to grasp the reason for their disagreement, leading to incorrect stance interpretations.
To catalyze progress beyond these limitations, we therefore introduce U-MStance, the first dataset designed for user-centric multimodal conversational stance detection. Comprising over 40k annotated comments across six real-world targets, U-MStance uniquely incorporates multimodality in both source posts and conversational replies, providing a realistic and challenging benchmark that captures the nuances of user-centric, multimodal interactions. Building upon this critical new dataset, we further propose PRISM, a Persona-Reasoned multImodal Stance Model for MCSD. Specifically, PRISM first explores the underlying personal tendencies behind stance expressions and derives longitudinal user persona traits inspired by Big Five (OCEAN) theory [14] from historical posts and comments. Then, it introduces a rationalized cross-modal grounding module, leveraging Chain-of-Thought (CoT) reasoning to explicitly bridge the profound semantic and pragmatic gap between modalities in conversational contexts. Ultimately, these personas and contextual representations are fed into a mutual task reinforcement mechanism, which jointly optimizes stance classification and stance-aware response generation, enabling bidirectional knowledge transfer that enhances both detection performance and the modeling of user communication styles. Comprehensive experiments on our U-MStance dataset show that PRISM achieves significant improvements over strong baselines across multiple metrics. The main contributions can be summarized as follows:
@@ 2.1 Stance Datasets
The development of stance detection datasets has evolved from analyzing isolated texts to modeling complex interactions. Early benchmarks like SemEval-2016 Task 6 [30] and VAST [3] focused on sentence-level textual stance detection, establishing foundational models but overlooking the context of dialogue. This limitation spurred the creation of conversational datasets such as SRQ [36] and CANT-CSD [22], which treated stance as a dynamic process unfolding within a conversation. The most recent advancement is the shift toward multimodal stance detection, which incorporates visual elements alongside text to better reflect modern social media discourse. However, even the latest datasets for MCSD, such as MmMtCSD [31], still face two key shortcomings that hinder realistic stance modeling. First, multimodality is constrained to source posts as static background context, while conversational replies are restricted to text-only. In addition, information about users, including their posting history, is largely ignored, overlooking contextual factors that shape an individual’s expressed stance and leaving a substantial gap in building comprehensive and authentic stance detection models.
@@ 2.2 Stance Detection Methods
Methodological advancements in stance detection has spanned multiple task settings in-target [19], cross-target [11, 26, 40], and zero-shot scenarios [2, 24, 46]. Early deep learning approaches primarily fine-tuned pre-trained models such as BERT [5]. However, the field was revolutionized by the emergence of Large Language Models (LLMs), whose powerful semantic understanding and adaptability introduced new capabilities [6, 7, 18, 12].
Following this paradigm shift, LLM-centric research has moved beyond direct prompting or CoT reasoning toward increasingly sophisticated methodologies [45, 44]. These include fine-tuning with external knowledge [17], prompt-based multi-perspective reasoning [11, 37], formalizing logic structures [43], and designing multi-agent collaborative systems [16, 38]. Despite this rapid progress, current approaches often treat users as homogeneous, overlooking the personal traits and discourse tendencies that shape stance expression. This limits the interpretation of personalized reactions and conflicting views, necessitating a shift from content-centered to user-centric modeling.
@@ 3 Dataset
This section outlines the construction and key attributes of our U-MStance dataset. As illustrated in Figure 2, our data construction pipeline includes collection, processing and verification to ensure data quality and relevance. Subsequently, data analysis is conducted to profile the dataset’s composition and validate its quality.
@@ 3.1 Data Construction
Dataset Collection. To capture realistic user interactions, we sourced conversational data from Reddit [^1] https://www.reddit.com, a large and diverse online forum, using the PRAW library. Covering the period from 2017 to 2025, we initially identified 835 posts related to six high-engagement, real-world targets: Trump, Biden, Tesla, BMW, Costco, and Bitcoin. These posts were selected based on popularity metrics like upvotes and comment counts to ensure a sufficient volume of conversational replies. We also collected each user’s historical posts and comments to enrich their personal information.
@@
Pre-annotation. Following this, GPT-4o-mini was employed for automated pre-annotation, assigning a preliminary stance label (Favor, Against, or None) to each instance relative to its target.
Human Verification. The automated labels underwent a rigorous human verification process by seven NLP researchers. Annotators were first qualified through two pilot rounds and an expert review. During the main annotation, each instance was independently reviewed by at least two annotators. Disagreements were resolved by a third senior annotator, and the final labels were determined through majority voting. Beyond stance correction, annotators also evaluated the contextual relevance of any associated images. This multi-annotator workflow ensures high label fidelity, and reinforces the overall reliability of the dataset.
Table 1: Comparison of U-MStance with existing datasets. Size: the number of annotated instances; Target: the number of the targets; Conv.: support for the conversational context; Vision-Related: percentage of vision-integrated instances; User: support for user information; -: not applicable.
Dataset Size Target Conv. Vision-Related User
SEM16 [30] 4,870 6
P-stance [20] 21,574 3
COVID-19 [13] 6,133 4
WT-WT [9] 51,284 5
VAST [3] 23,525 -
SRQ [36] 1,348 4
CANT-CSD [22] 5,376 1
CTSDT [23] 53,861 1
MT-CSD [33] 15,876 5
C-MTCSD [34] 24,264 5
MT2-CSD [32] 24,457 6
MMVax [41] 11,300 - 100%
MMSD [27] 17,544 5 100%
MmMtCSD [31] 21,340 3 65.99%
U-MStance(Ours) 40,003 6 100%
Figure 3: Overview of our PRISM framework.
@@ 3.2 Data Analysis
Data Statistics. Our dataset comprises 40,003 annotated instances and 24952 user profiles. Each instance consists of a complete conversational thread accompanied by its associated images within the conversation. Crucially, Table 1 provide a comparison between U-MStance and existing datasets. U-MStance is the first dataset in this domain to systematically incorporate user information, offering a foundational resource to address the prevalent issue of user homogeneity. More comprehensive statistics, including the distributions of labels and conversational depths, are detailed in Appendix A. For model development, the dataset is partitioned into training, validation, and test sets using an 8:1:1 ratio. We strictly partition the data by cross-referencing timestamps to ensure that the three sets are mutually independent.
Quality Assessment. To ensure the annotation reliability, we measured inter-annotator agreement using Cohen’s Kappa [8], achieving an average score of 0.64, indicative of substantial agreement. This consistency was further reinforced through a rigorous post-annotation review, in which senior researchers audited a randomly selected subset of the annotated data to verify accuracy and resolve borderline cases, providing an additional layer of quality assurance.
@@ 4.1 Longitudinal User Persona Distillation
Stance is an externalized reflection of a user’s underlying personal tendencies. Thus, each user is modeled as an entity whose stance expression is shaped by stable personality traits. For this purpose, we adopt the empirically validated Big Five (OCEAN) personality model [14], which characterizes personality along five fundamental dimensions: Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism. To infer these traits for user $u_{N}$, we first aggregate their complete history $H_{u_{N}}$, which comprises all historical posts and comments.
@@
where each rating $r\in\{1,2,3,4,5\}$. The resulting profile $p_{u_{N}}$ is a concise, structured text string that serves as a crucial, personalized context vector. This profile is explicitly injected into the prompt for the final stance detection task, ensuring the model’s prediction is conditioned on the stable identity of the user in question. The prompt template can be found in Table 7 of Appendix C.
@@ 4.2 Rationalized Cross-Modal Grounding
Images in online conversations are not decorations but communicative acts whose pragmatic meaning is shaped by the conversational context. To decipher this context-dependent rhetorical intent, we propose the Rationalized Cross-Modal Grounding (RCMG) module to generate an intent-aware caption for each image in the conversation. Inspired by Chain-of-Thought [39], this module employs a two-stage reasoning process to bridge the semantic-pragmatic gap. The complete prompt template can be found in Table 8 of Appendix C.
@@
The resulting caption $\hat{x}_{i}$ captures the rhetorical function of the image within the conversation.
@@ 4.3 Mutual Task Reinforcement
To cultivate a comprehensive, user-centric understanding of stance, PRISM employs a multitask learning paradigm that jointly optimizes two inherently correlated tasks. The core intuition is that by training the model to generate a user’s response as an auxiliary task, it derives a deeper understanding of the pragmatic and user-specific cues essential for the primary task of stance prediction. This mutual reinforcement serves as a robust regularizer, enhancing the stability and prediction accuracy.
@@
The classification loss is the negative log-probability of the ground-truth label $y_{N}$.
@@ 5.1 Experimental Setup
Implementation details. Our framework is built on Qwen2.5-VL-7B [4] and fine-tuned for 5 epochs using our multitask objective, with the loss-weighting hyperparameter $\lambda$ set to 0.7. We use the AdamW optimizer [1], with an initial learning rate of $1\times 10^{-5}$, a cosine decay schedule, and a global batch size of 16. All experiments were run on a server with four NVIDIA A100 (40 GB) GPUs using PyTorch and the Hugging Face Transformers library.
Evaluation metrics. We adopt the average F1 score (F1-avg) as the primary evaluation metric, following [31] and [34]. F1-avg is computed as the mean of F1 scores for the against and favor stances, denoted as F1-against and F1-favor.
@@ 5.2 Baseline Methods
We conduct extensive experiments with a range of existing methods, categorized into text-only and multimodal approaches. All baseline models are fine-tuned on our specialized training data, with the exception of the GPT series, which is evaluated in a zero-shot setting. Text-only baselines. We evaluate the following methods: BERT [10], RoBERTa [29], PoliBERTweet [15], LLaMA2 [35], GPT-3.5 and GPT-4. Multimodal baselines. We include: BERT-ViT [27], TMPT [27], LLaVA [28], Qwen-VL [4], MiMo [42] and GPT4-1. For more details, please refer to Appendix B.